Given the function f(x) = −2x^2x + 5x − 1, and g(x) = 5x + 3 evaluate the following:
a. f(−2) b.f(3) + 7 c. g(x − 1)

Answer by Boreal(15235)

f(-2)--put -2 everywhere you see an x in the first function
-2(-2)^(2*-2)+5(-2)-1=-2^-4 + 5(-2)-1=-2^(-4)-11=-1/16-11=-11 (1/16)
-2(6)+15+7=-64=22=-42
=5(x-1)+3=5x-2